### v1

# LlamaLens: Specialized Multilingual LLM for Analyzing News and Social Media Content

Abstract
[…]
LlamaLens: Specialized Multilingual LLM for
Analyzing News and Social Media Content
[…]
LlamaLens: Specialized Multilingual LLM for
Analyzing News and Social Media Content
[…]
1 Introduction
Figure 1: Capabilities, tasks and associated datasets in LlamaLens.
LLMs have significantly advanced the field of AI, demonstrating capabilities in solving downstream NLP tasks and exhibiting knowledge understanding and cognitive abilities Touvron et al. (2023); Mousi et al. (2024). However, extending these capabilities with more domain-specific knowledge and achieving higher accuracy requires domain specialization. This entails customizing general-purpose LLMs with domain-specific data, augmented by relevant domain knowledge Ling et al. (2023). One prominent area where LLMs can be customized with specialized knowledge is the news and social media analysis. Since their emergence, the use of AI in news production, analysis, video scripting, copyediting, translation, and dissemination has grown significantly Shi and Sun (2024); Simon (2024). In addition to news production and dissemination, a closely related area is social media content analysis Zeng et al. (2024); Liu et al. (2024); Franco et al. (2023); Hasanain et al. (2024b). This growing range of applications creates a strong demand for specialized LLMs to support journalists, fact-checkers, communication specialists and social media users. There has been an attempt to develop a tool [^2] https://newsgpt.ai/ based on ChatGPT to support journalists in news production and delivery Hireche et al. (2023); however, it relies solely on ChatGPT for language understanding and response generation in response to reporters’ queries. Other efforts to support journalists include tools for creating news reels Wang et al. (2024c), classifying news frames using GPT Alonso del Barrio and Gatica-Perez (2023), generating image captions for news articles Anagnostopoulou et al. (2024), news recommendation systems Wang et al. (2024b), and specialized models for the news media and business sectors Bao et al. (2024). However, little to no attention has been given to developing specialized models for news and social media content analysis (see related work in Section 2.3). Such models are crucial for tasks like identifying whether a news article or social media post contains a claim, assessing its factual accuracy, determining its relevance for fact-checking, and evaluating whether the content is offensive, incites hate, or requires moderation. To address this gap, we focused on developing specialized LLMs by fine-tuning an existing model to better support journalists, fact-checkers, communication specialists, and social media analysts. As illustrated in Figure 1, our goal was to equip LLMs with a range of specialized capabilities across multiple languages. Our main contributions are summarized below.
  • We develop and release a specialized LLM, LlamaLens, which covers 5 broader capabilities, associated with 19 tasks and 52 datasets in three languages: Arabic, English, and Hindi.
  • We develop and release an instruction-following dataset, developed using a semi-supervised approach.
  • We explore various data shuffling techniques, based on language, dataset, and task, during training and present our findings.
  • We present detailed experimental results comparing with (i) Llama-3.1-8B-Instruct model, (ii) quantized models trained over data with different shuffling, and (iii) state-of-the-art baselines using dataset-specific metrics.
Our findings suggest that (i) LlamaLens acquires domain- and language-specific knowledge, highlighting the importance of specialized models, (ii) considerably smaller versions of the model (resulting from fine-tuning with quantization) acquire that knowledge in comparison to the un-finetuned model, showing significantly better performance, and (iii) comparison to the SOTA suggests that there is still a room for improvement.
2 Related Work
2.1 LLMs for Journalism
Recent studies have explored the intersection of LLMs, journalism, and social media, shedding light on both the opportunities and challenges of integrating AI into news reporting Cheng (2024); Petridis et al. (2023); Hasan et al. (2024); Quinonez and Meij (2024); Nishal and Diakopoulos (2024); Ding et al. (2023). For example, Brigham et al. (2024) and  Breazu and Katsos (2024) examined the use of LLMs like GPT-4 in journalistic workflows, focusing on the ethical, and quality implications and generating narratives. LLMs have also been integrated in news production, focusing on its benefits and ethical challenges Shi and Sun (2024). One key challenge of using generic LLMs in journalism is their tendency to generate false or misleading information Cheng (2024); Augenstein et al. (2023), a phenomenon known as hallucination. Nishal and Diakopoulos (2024) also highlighted key concerns, including hallucinations, factual inaccuracies, and the potential threat to journalistic objectivity. To address these issues, they proposed a value-sensitive design approach, advocating for AI systems that offer transparent explanations, explicitly represent uncertainty, and give journalists more control over the generated content. Bloomberg has integrated LLMs into its news production processes Quinonez and Meij (2024), aiming to enhance automation while preserving essential journalistic principles such as accuracy and transparency. Similarly,  Ding et al. (2023) examined the role of LLMs in human-AI collaboration, particularly for generating news headlines. To tackle content creation challenges on visual platforms like Instagram Reels and TikTok, ReelFramer, a multimodal writing assistant, was developed Nickerson et al. (2023). Additionally,  Cheng (2024) emphasized the need for customized LLMs tailored to news reporting, proposing solutions like supervised fine-tuning and constitutional AI, which integrates reinforcement learning from AI feedback to combat misinformation and rebuild reader trust. To facilitate science journalism,  Jiang et al. (2024) introduced a novel approach that leverages collaboration among multiple LLMs to improve the readability and clarity of news articles.
2.2 News and Social Media Analysis
For news and social media analysis there has been research effort with a special focus on fact-checking and harmful content detection Lee et al. (2024); Quelle and Bovet (2024), and news reliability classification Ibrahim (2024). Quelle and Bovet (2024) demonstrated that LLM agents can be employed for fact-checking by retrieving relevant evidence and verifying claims. Similarly,  Ibrahim (2024) explored fine-tuned LLMs, such as Llama-3, to automate the classification of reliable versus unreliable news articles, specifically in Spanish. The Enhancing Perception Hsu et al. (2024) and FACT-GPT Choi and Ferrara (2024) frameworks tackle misinformation by refining fake news explanations through a conversational refinement approach. Similarly, VerMouth Russo et al. (2023) automates social media fact-checking, contributing to broader efforts in combating misinformation. Additionally, the expert recommendation framework Zhang et al. (2024b) leverages a multi-layer ranking system with LLMs, balancing reliability, diversity, and comprehensiveness when suggesting experts for news events. Other initiatives include Botlitica Musi et al. (2024), which identifies propagandistic content in political social media posts, and JSDRV Yang et al. (2024), which focuses on stance detection and rumor verification. In the realm of investigative journalism,  Ali (2024) introduced a tool to retrieve and summarize relevant documents, while  Alonso del Barrio et al. (2024) focused on detecting framing in television program content. Addressing political bias,  Trhlik and Stenetorp (2024) explored bias identification using LLMs. Additionally,  Alonso del Barrio et al. (2024) proposed prompt-engineering LLMs to analyze framing in spoken content from television programs. A comprehensive study was conducted by Zeng et al. (2024), highlighting the use of LLMs in social media applications.
2.3 Specialized LLMs
Ling et al. (2023) highlighted the importance of developing specialized models for several reasons. One key reason is that, much like humans, acquiring domain expertise and capabilities often requires years of training and experience. Therefore, it is important to train LLMs with domain knowledge to serve professional level usage. In this direction, a recent work by Kotitsas et al. (2024) explored fine-tuning LLMs to improve claim detection. Bao et al. (2024) trained an LLM, FLLM, using curated knowledge with a focus on the business and media domains. For training, they utilized articles published by Fortune Media. The OpenFactCheck framework Wang et al. (2024d) tackles the evaluation of factual accuracy in LLM-generated content. This customizable architecture enables the assessment of both LLM factuality and fact-checking systems, promoting standardized evaluations essential for advancing research on the reliability and factual correctness of LLMs.  Wang et al. (2024a) proposed an explainable fake news detection framework that uses LLMs to generate and evaluate justifications from opposing parties. A defense-based inference module then determines veracity, improving detection accuracy and justification quality, as demonstrated on two benchmarks. In contrast to prior work, our research focuses on developing a specialized model for a wide range of tasks and capabilities for news and social media analysis, representing the first effort in this direction with multilingual capabilities.
3 Tasks and Datasets
3.1 Data Curation
For dataset curation, we selected key capabilities and their associated tasks, as illustrated in Figure 1, and identified publicly available datasets aligned with these tasks. The choice of languages was influenced by their prominence in the Arabian Peninsula region, making them essential for analyzing linguistic and cultural diversity in the context of social and news media in the region. The initial collection consists of 103 datasets, some of which we excluded due to their different versions (e.g., we selected ArSarcasm-v2 Abu Farha et al. (2021) instead of version 1). After initial pre-selection, the resulting collection consists of 52 datasets as detailed in Tables 3, 4 and 5 in Appendix B. The datasets span various sources, including social media posts, news articles, political debates and transcripts. It consists of $\sim$2.7m samples and a total of 222 labels, reflecting the complexity of tasks such as checkworthiness, claim detection, cyberbullying, emotion detection, news categorization, and more.
3.2 Preprocessing
After collecting the datasets, we observed that while most were pre-divided into training, development, and test sets, 18 datasets lacked these splits. In cases where datasets were not pre-split, we partitioned them into 70% for training, 20% for testing, and 10% for development. For datasets containing only training and test sets, we further divided the training set, allocating 30% for development. To preserve the class distribution across splits, we employed stratified sampling. We applied several other preprocessing steps such as removing duplicates, unifying labels (e.g., check-worthiness to checkworthiness, fixing uppercase to lowercase), and removed entries with less than 3 letters. These preprocessing steps ensured that the datasets were clean, well-structured, and ready for subsequent analysis or model training. After preprocessing, we obtained a total of 1.2m, 1.4m and 0.14m samples for Arabic, English, and Hindi, respectively. The number of labels in the datasets ranges from 2 to 42. The datasets also include both multiclass and multilabel setups. We provide distribution of the datasets, number of labels and their splits in Tables 3, 4 and 5 in Appendix B. The datasets come in different sizes, ranging from the smallest (e.g., CT-24 subjectivity) to the largest (e.g., English news summarization dataset), and with varying label distributions, from skewed (e.g., propaganda) to more balanced (e.g., Arabic CT22 claim detection). In Appendix A, we provide detailed descriptions of the tasks and datasets.
4 Methodology
Figure 2: Approach for the LlamaLens: datasets, model training, and evaluation.
In Figure 2, we present the methodological steps involved in the development of LlamaLens, which is also formalized in Algorithm 1. In Section 3, we discussed the details of the dataset curation and preprocessing. The following subsections discuss the remaining steps in the development process.
4.1 Instruction Dataset
Our approach follows the typical and effective pipeline of aligning LLMs with user intentions and tasks by LLM fine-tuning on representative data. Such approach usually involves creating instruction datasets starting from existing NLP datasets Longpre et al. (2023). An instruction sample is a triplet of a natural language instruction describing the task, an optional input, and an output that represents the answer following the instruction Wang et al. (2023).
Natural language instructions
There are several potential techniques to create the needed natural language instructions, including manual and automatic approaches. As instructions diversity positively affects model performance and generalization Dubois et al. (2024); Pang et al. (2024); Zhang et al. (2024a), we aim to create a diverse instruction dataset. Due to the scale of tasks and datasets of focus in this work, creating such a diverse set manually is time-consuming and can lead to limited instruction styles Wang et al. (2022); Kaur et al. (2024). We opt to automatically generate instructions by prompting two highly-effective closed LLMs, GPT-4o  [^3] https://platform.openai.com/docs/models/gpt-4o and Claude-3.5 Sonnet,  [^4] https://www.anthropic.com/news/claude-3-5-sonnet to generate a diverse set of 10555We chose the number 10 as a compromise between diversity and the number of samples per instruction. Finding an optimal value requires further study. English instructions per LLM, resulting in 20 instructions per dataset. To ensure the models generate instructions fitting our datasets, we provide the models with the datasets metadata, including dataset name, language, task, task definition and labels space. Exact prompt used to generate instructions and examples of generated instructions can be found in Appendix C. It is important to note that we only require English instructions, as a previous study demonstrated that English prompts outperform language-specific native prompts Kmainasi et al. (2024). Finally, for each input dataset of the 52, we create instructions by appending a randomly selected natural instruction, of the generated 20, per example of each training subset. This guarantees versatile instruction styles even for the same input dataset. Our final instruction tuning dataset is the set of the prepared instructions for all datasets.
Input: Set of datasets $\mathcal{D}=\{D_{1},D_{2},\dots,D_{N}\}$
Input: $LLM_{c}$, $LLM_{g}$, $LLM_{f}$
Output: Fine-Tuned Model $LLM_{f}^{\prime}$
1
[…]
3
4for $k\leftarrow 1$ to $N$ do
5       $I_{k}=\operatorname{GenerateInstruct}(LLM_{c},LLM_{g},D_{k})$
6      foreach data point $d\in D_{k}$ do
7             $i\sim\text{Uniform}(I_{k})$
8             $d^{\prime}=(d,i)$
9             $D^{\prime}\leftarrow D^{\prime}\cup\{d^{\prime}\}$
10       end foreach
11
12 end for
13Fine-tune the model $LLM_{f}$ using $D^{\prime}$: $LLM_{f}^{\prime}=\operatorname{FineTune}(LLM_{f},D^{\prime})$ return Fine-Tuned Model $LLM_{f}^{\prime}$;
Algorithm 1 Algorithm for the dataset creation and LlamaLens model development. $LLM_{c}$ and $LLM_{g}$ refers to Claude 3.5 and GPT-4o models, respectively. $D^{\prime}$ is the final instruction dataset.
4.2 Training
Our experiments are based on the most effective open LLM to-date, Llama 3.1, which shows effective performance even in multilingual setups Dubey et al. (2024). Fine-tuning larger scales of the model (e.g., 70B version) holds a great overhead in terms of time and computational cost. Moreover, these models might be inaccessible to the greater research community. Thus, we opt to focus on the smallest Llama 3.1-8B version. We particularly base our LlamaLens model on the instruct version of Llama 3.1-8B as it is already aligned to several user tasks.
4.2.1 Training Setups
We instruction-tune Llama 3.1-8B-Instruct [^6] We use the term Llama-instruct to refer to this model in the rest of the paper. following different setups. Given that fine-tuning LLMs typically requires substantial computational resources, making it a time-consuming and resource-intensive process, therefore, to address this challenge, our main LlamaLens model is fine-tuned in full precision (16-bit) following parameter-efficient fine-tuning using Low-Rank Adaptation (LoRA) Hu et al. (2021). In addition to the full precision model, we aimed to train smaller models, achieving two goals: (i) release smaller but effective models to be used in resource-constrained environments, and (ii) efficiently investigate the effects of some parameter settings on model performance, to guide the full model training. Thus, we also fine-tune the original Llama-8B-Instruct model employing QLoRA Dettmers et al. (2024), which involves quantization of the model’s weights and significantly enhances memory optimization, while maintaining acceptable performance.
Task Dataset Metric SOTA Llama Ours $\Delta$ Task Dataset Metric SOTA Llama Ours $\Delta$
Arabic English
Att.worthiness CT22Attentionworthy W-F1 0.412 0.158 0.544 0.132 Checkworthiness CT24_T1 F1_Pos 0.753 0.404 0.877 0.124
Checkworthiness CT24_T1 F1_Pos 0.569 0.404 0.877 0.308 Claim claim-detection Mi-F1 0.545 0.915
Claim CT22Claim Acc 0.703 0.581 0.778 0.075 Cyberbullying Cyberbullying Acc 0.907 0.175 0.847 -0.060
Cyberbullying ArCyc_CB Acc 0.863 0.766 0.753 -0.110 Emotion emotion Ma-F1 0.790 0.353 0.878 0.088
Emotion Emotional-Tone W-F1 0.658 0.358 0.748 0.090 Factuality News_dataset Acc 0.920 0.654 0.946 0.026
Emotion NewsHeadline Acc 1.000 0.406 0.551 -0.449 Factuality Politifact W-F1 0.490 0.121 0.290 -0.200
Factuality Arafacts Mi-F1 0.850 0.210 0.534 -0.316 News Cat. CNN_News_Articles Acc 0.940 0.644 0.915 -0.025
Factuality COVID19Factuality W-F1 0.831 0.492 0.781 -0.050 News Cat. News_Category Ma-F1 0.769 0.970 0.505 -0.264
Harmfulness CT22Harmful F1_Pos 0.557 0.507 0.508 -0.049 News Genre SemEval23T3-ST1 Mi-F1 0.815 0.687 0.241 -0.574
Hate Speech annotated-hatetweets-4 W-F1 0.630 0.257 0.549 -0.081 News Sum. xlsum R-2 0.152 0.074 0.141 -0.010
Hate Speech OSACT4SubtaskB Mi-F1 0.950 0.819 0.802 -0.148 Offensive Offensive_Hateful Mi-F1 0.692 0.805
News Cat. ASND Ma-F1 0.770 0.587 0.938 0.168 Offensive offensive_language Mi-F1 0.994 0.646 0.884 -0.110
News Cat. SANADAkhbarona Acc 0.940 0.784 0.922 -0.018 Offensive & Hate hate-offensive-speech Acc 0.945 0.602 0.924 -0.021
News Cat. SANADAlArabiya Acc 0.974 0.893 0.986 0.012 Propaganda QProp Ma-F1 0.667 0.759 0.851 0.184
News Cat. SANADAlkhaleej Acc 0.986 0.865 0.967 -0.019 Sarcasm News-Headlines Acc 0.897 0.668 0.956 0.059
News Cat. UltimateDataset Ma-F1 0.970 0.376 0.883 -0.087 Sentiment NewsMTSC Ma-F1 0.817 0.628 0.627 -0.190
News Credibility NewsCredibility Acc 0.899 0.455 0.494 -0.405 Subjectivity CT24_T2 Ma-F1 0.744 0.535 0.508 -0.236
News Sum. xlsum R-2 0.137 0.034 0.075 -0.063 Hindi
Offensive ArCyc_OFF Ma-F1 0.878 0.489 0.652 -0.226 Factuality fake-news Mi-F1 0.759 0.713
Offensive OSACT4SubtaskA Ma-F1 0.905 0.782 0.899 -0.006 Hate Speech hate-speech-detection Mi-F1 0.639 0.750 0.994 0.355
Propaganda ArPro Mi-F1 0.767 0.597 0.762 -0.005 Hate Speech Hindi-Hostility W-F1 0.841 0.469 0.720 -0.121
Sarcasm ArSarcasm-v2 F1_Pos 0.584 0.477 0.307 -0.277 NLI NLI_dataset W-F1 0.646 0.633 0.655 0.009
Sentiment ar_reviews_100k F1_Pos 0.343 0.665 News Sum. xlsum R-2 0.136 0.078 0.117 -0.019
Sentiment ArSAS Acc 0.920 0.603 0.795 -0.125 Offensive Offensive Speech Mi-F1 0.723 0.621 0.847 0.124
Stance stance Ma-F1 0.767 0.608 0.936 0.169 Cyberbullying MC-Hinglish1.0 Acc 0.609 0.233 0.587 -0.022
Stance Mawqif-Arabic-Stance Ma-F1 0.789 0.764 0.867 0.079 Sentiment Sentiment Analysis Acc 0.697 0.552 0.669 -0.028
Subjectivity ThatiAR F1_Pos 0.800 0.562 0.207 -0.593
Table 1: LlamaLens performance across all datasets. SOTA: State-of-the-art results. Llama: Llama 3.1-8B-Instruct. R-2: ROUGE-2, Acc: Accuracy, Mi-F1: Micro-averaged F1, Ma-F1: Macro-averaged F1, W-F1: weighted F1, F1_Pos: F1 of the positive class, –: No SOTA scores found. NLI: Natural Language Inference. Att.worthiness: Attentionworthiness. News Sum.: News Summarization. Cat.: Categorization. : Data was not pre-split.
4.2.2 Experimental Setup
Datasets Sampling
[…]
Dataset Shuffling
The order of instructions in the training dataset can significantly impact model performance. For example, Pang et al. (2024) demonstrate that ordering instructions by complexity influences the effectiveness of tuned models. In light of this, we investigate how different orders of samples affect the performance of LlamaLens, employing four different data shuffling and ordering techniques to identify the optimal sequence. Although there are several possible ordering configurations, however, we limit the scope of this study to the effects of language and task order.
  1. 1.
    Alphabetically ordered: This is a basic setup where languages and datasets are ordered alphabetically—Arabic, followed by English, and then Hindi—without shuffling.
  2. 2.
    Shuffled by language: Randomly shuffle datasets while preserving the order of languages.
  3. 3.
    Shuffled by task: Tasks are organized alphabetically, with datasets shuffled across tasks regardless of language.
  4. 4.
    Fully randomized: Complete randomization of the training dataset points.
Parameters Setup
For all models we train, we set a LoRA learning rate of 2e-4 alongside a linear learning rate schedule. Optimization was performed using AdamW Loshchilov and Hutter (2017), with a batch size of 16. All experiments were executed on four NVIDIA H100-80GB GPUs. In the first set of experiments, we trained four models quantized to 4-bit precision using QLoRA. Although the models store weights in 4-bit format, computations are performed in BFLOAT16 (bf16), with both the LoRA rank and $\alpha$ set to 16. Each model was trained on one of the dataset shuffling configurations. After identifying the optimal order of the training dataset—based on average model performance on our test sets, as shown in Figure 3—we used that dataset order to fine-tune our LlamaLens model in full precision (16-bit). Due to the scale of the model and training set size, we train the model for two epochs, increasing LoRA rank to 512, following the recommendations in Xin et al. (2024), which suggests that higher ranks yield improved performance for multitask learning. LoRA $\alpha$ was set equal to the rank.
4.3 Evaluation
4.4 Evaluation Process
For evaluation, we employed a zero-shot approach, in which we directly prompted the models (original and fine-tuned) to perform tasks from the testing sets. The employed natural instruction/prompt is the first generated instruction (Section 4.1) for each dataset. To ensure reproducibility, we set the top_p to 1 and the temperature to 0. For classification tasks, we restrict the models to generate a maximum of 20 new tokens. For summarization, we applied a stopping criterion based on encountering a period (‘.’) or the end-of-sequence (eos_token).
4.5 Post-processing
As models can generate text beyond that required in the instruction, a post-processing method was implemented to extract labels from the generated models’ responses. Initially, a regular expression was used to accurately identify and extract the labels. Several transformations were also applied, including converting all text to lowercase, removing special characters, and addressing code-switching by replacing non-Latin characters with their Latin equivalents. A similar post-processing approach was reported by Abdelali et al. (2024).
4.6 Evaluation Metrics
All models were evaluated using standard classification metrics including weighted-, macro-, micro- F1 and accuracy. We use ROUGE-2 to evaluate the summarization task. Specifically, we use the same metric reported in state-of-the-art (SOTA) per dataset.
Dataset Metric Llama Alpha. Full Task Lang. Dataset Metric Llama Alpha. Full Task Lang.
Arabic English
CT22Attentionworthy W-F1 0.158 0.281 0.299 0.293 0.340 CT24_T1 F1_Pos 0.404 0.538 0.583 0.893 0.288
CT24_T1 F1_Pos 0.404 0.538 0.583 0.893 0.288 claim-detection Mi-F1 0.545 0.895 0.891 0.884 0.898
ans-claim Ma-F1 0.335 0.427 0.528 0.454 0.473 Cyberbullying Acc 0.175 0.664 0.794 0.781 0.764
CT22Claim Acc 0.581 0.712 0.735 0.715 0.723 emotion Ma-F1 0.353 0.647 0.662 0.584 0.654
ArCyc_CB Acc 0.766 0.767 0.818 0.840 0.776 News_dataset Acc 0.654 0.502 0.712 0.787 0.614
Emotional-Tone W-F1 0.358 0.595 0.635 0.641 0.609 Politifact W-F1 0.121 0.210 0.241 0.252 0.262
NewsHeadline Acc 0.406 0.316 0.319 0.387 0.288 CNN_News_Articles Acc 0.644 0.897 0.919 0.904 0.911
Arafacts Mi-F1 0.210 0.376 0.263 0.466 0.362 News_Category Ma-F1 0.970 0.964 0.913 0.635 0.668
COVID19Factuality W-F1 0.492 0.794 0.733 0.595 0.780 SemEval23T3-ST1 Mi-F1 0.687 0.325 0.494 0.470 0.590
CT22Harmful F1_pos 0.507 0.539 0.565 0.473 0.539 xlsum R-2 0.074 0.088 0.126 0.116 0.101
annotated-hatetweets-4 W-F1 0.257 0.436 0.311 0.371 0.394 Offensive_Hateful Mi-F1 0.692 0.791 0.768 0.792 0.778
OSACT4SubtaskB Mi-F1 0.819 0.946 0.901 0.910 0.911 offensive_language Mi-F1 0.646 0.893 0.871 0.657 0.821
ASND Ma-F1 0.587 0.790 0.787 0.803 0.811 hate-offensive-speech Acc 0.602 0.874 0.901 0.909 0.903
SANADAkhbarona Acc 0.784 0.924 0.904 0.930 0.938 QProp Ma-F1 0.759 0.773 0.803 0.751 0.773
SANADAlArabiya Acc 0.893 0.975 0.973 0.973 0.973 News-Headlines Acc 0.668 0.959 0.961 0.953 0.960
SANADAlkhaleej Acc 0.865 0.929 0.920 0.916 0.929 NewsMTSC-dataset Ma-F1 0.628 0.640 0.669 0.685 0.613
UltimateDataset Ma-F1 0.376 0.742 0.594 0.647 0.673 CT24_T2 Ma-F1 0.535 0.464 0.440 0.554 0.379
NewsCredibility Acc 0.455 0.665 0.845 0.904 0.600 Hindi
xlsum R-2 0.034 0.058 0.058 0.061 0.063 fake-news Mi-F1 0.759 0.802 0.633 0.567 0.653
ArCyc_OFF Ma-F1 0.489 0.835 0.846 0.856 0.836 hate-speech-detection Mi-F1 0.750 0.910 0.898 0.903 0.879
OSACT4SubtaskA Ma-F1 0.782 0.876 0.852 0.863 0.849 Hindi-Hostility W-F1 0.469 0.666 0.564 0.664 0.526
ArPro Mi-F1 0.597 0.660 0.623 0.696 0.655 NLI_dataset W-F1 0.633 0.516 0.573 0.537 0.564
ArSarcasm-v2 F1_Pos 0.477 0.154 0.542 0.429 0.472 xlsum R-2 0.078 0.074 0.094 0.095 0.080
ar_reviews_100k F1_Pos 0.343 0.552 0.626 0.614 0.626 Offensive_Speech Mi-F1 0.621 0.692 0.701 0.733 0.763
ArSAS Acc 0.603 0.780 0.774 0.763 0.772 MC-Hinglish1.0 Acc 0.233 0.438 0.435 0.560 0.545
stance Ma-F1 0.608 0.634 0.774 0.775 0.853 Sentiment Analysis Acc 0.552 0.640 0.643 0.636 0.624
Mawqif-Arabic-Stance Ma-F1 0.764 0.774 0.819 0.845 0.846
ThatiAR F1_Pos 0.562 0.558 0.544 0.574 0.591
Table 2: Quantized models performance across all datasets and dataset shuffling techniques. R-2: ROUGE-2. Acc: Accuracy. Mi-F1: Micro-averaged F1. Ma-F1: Macro-averaged F1. W-F1: weighted F1. F1_Pos: F1 of the positive class. Llama: Llama 3.1-8B-Instruct. Alpha.: Quantized model trained over dataset ordered alphabetically. Full: Quantized model trained over fully randomized dataset. Task: Quantized model trained over dataset shuffled by task. Lang.: Quantized model trained over dataset shuffled by language. Numbers in bold represent the best performance per dataset.
5 Results and Discussions
5.1 Data Shuffling Effect
In Figure 3, we present the results as averages computed across datasets and languages. Detailed results are provided in Table 2. Among the four different ordering strategies, shuffling by task outperformed other configurations. Both shuffling by language and alphabetic ordering showed comparable performance. Based on this experiment, we elect to train our LlamaLens model following the shuffling by task data ordering approach.
Figure 3: Impact of data shuffling technique on fine-tuned quantized Llama-3.1 performance.
5.2 Results on LlamaLens
In Table 1, we report the full results of our LlamaLens model across the different languages. Overall, LlamaLens significantly outperforms the Llama-instruct with an average improvement of 33%. Language wise LlamaLens outperforms Llama-instruct on average 32.5%, 32.2% and 29.5% for Arabic, English and Hindi, respectively. For Arabic, out of 26 tasks, LlamaLens only under-performs on cyberbullying, sarcasm and one of the hate-speech datasets. For English, out of 17 tasks LlamaLens outperforms on 14 tasks and lower performing tasks include sentiment, subjectivity and news categorization. For Hindi, in all 8 datasets LlamaLens outperforms Llama-instruct. In comparison to SOTA’s average performance of 0.75, LlamaLens average performance is 0.69. We should note here that since 18 out of the 52 datasets were not pre-split (Section 3.2), SOTA on these datasets is not directly comparable to our model, as the testing splits might be different. Computing average performance excluding these datasets reduces the gap, with a SOTA of 0.72 and LlamaLens performance of 0.67. In terms of number of datasets where LlamaLens improved over SOTA, we find that it outperforms SOTA in 16 test sets, and has a comparable performance (difference $\geq$ 0 and $<$-0.03) in 10 other testing sets. Dataset-specific improvements for English include tasks such as checkworthiness, emotion, news factuality, and propaganda. For Arabic, improvements have been observed in checkworthiness, emotion, and stance, while performance on news categorization is mixed, with improvements seen in 2 out of 5 datasets. For Hindi, significant improvements have been observed in hate speech, followed by NLI and offensive language.
5.3 Error Analysis of Llama-instruct Model
We conducted an error analysis on the responses of the Llama-instruct model, to identify challenges in handling text of low-resource languages on certain classification tasks. Several common issues emerged across these models, particularly in tasks related to offensive language, and hate-speech detection, factuality, and news categorization. One recurring problem was the inability of the model to provide labels in numerous instances, often responding with phrases like “I cannot provide a label” or “Arabic text is not easily classifiable into categories without context or translation.” Another observed issue is language confusion in output. Although the models were instructed to output labels exclusively in English, they occasionally returned labels in Arabic or code-switched responses, which is inline with language confusion reported in relevant studies Marchisio et al. (2024), however, differently, we showcase the confusion can occur at the smallest unit of a single character. For instance, in some cases, the model generated outputs like “\<ف>actual” (referring to “factual” where Arabic character \<ف> is a transliteration of character “f”), and “\<س>포츠” (Korean for “sport” transliterated as “seupocheu” where character \<س> is a transliteration of character “s”) despite no instructions involving Korean language. This highlights a phonetic-level code-switching phenomenon. It also occurred on longer sequences such as the model responding with “\<سار>castic” instead of “Sarcastic,” where “\<سار>” is actually pronounced similarly to “Sar”. In contrast, our fine-tuned versions of the model do not display such issues, with the models adhering more consistently to the instructed output languages.This suggests that fine-tuning is critical for improving language-specific performance.
6 Conclusion and Future Work
In this study, we propose a specialized model, LlamaLens, focused on news and social media analysis, designed to assist journalists, fact-checkers, and social media analysts. We curated 52 datasets spanning three prominent languages spoken in the Arabian Peninsula region: Arabic, English, and Hindi. To develop LlamaLens, we created an instruction-following dataset based on these curated datasets and fine-tuned the Llama 3.1-8B-Instruct model. Our experiments show that LlamaLens outperforms the SOTA on 16 datasets, performs comparably on 10 datasets, and underperforms on the rest of the datasets. However, on average, LlamaLens and its quantized versions significantly surpasses the Llama-instruct model. Our findings from error analysis suggests that it is important to inject specialized domain and language knowledge to obtain the desired outcome from Llama-instruct. Our future studies include experimenting with different rank orders and focusing on quantized version of the model to make it usable in a low-resource settings.
7 Limitations
Our experiments were focused on a single open LLM, further LLMs can be explored. The training datasets had a bigger representation of Arabic, but as experiments showed, the proposed model improved performance even on other languages. Further examination of the effect of training examples selection and shuffling techniques is needed to understand these effects on the model performance.
Ethics and Broader Impact
Our experiments were conducted on training datasets publicly released to the research community. We adhered to the licenses associated with them whenever available. Some of the data points we will be releasing as part of our instruction dataset contain vulgar, offensive, or disturbing content which is a natural occurrence on social media, thus caution is recommended for users of our dataset. The models and instruction dataset we release could be invaluable to news agencies, journalists, and social media platforms. However, we encourage developers and users of the models to be critical of their usage.
Appendix A Tasks and Datasets
In this section, we provide a comprehensive overview of the tasks and datasets used throughout our study. Each task is outlined with a brief description, followed by the associated datasets that were utilized. The datasets are presented based on the specific languages and their relevance to the tasks at hand, with a focus on the binary and multi-labeled classifications where applicable. The tasks cover a wide range of objectives, from detecting harmful content to classifying news articles, identifying emotions, and more. This appendix serves as a detailed reference to support the methodology and scope of the research.
A.1 Attentionworthiness
Attentionworthiness categorizes social media posts to determine whether they require attention and, if so, what kind of attention is needed. This task helps prioritize responses for policymakers by identifying critical posts that discuss actions, advice, or calls for intervention.
Dataset
For the Attentionworthiness Detection task, we utilized a subset of the Arabic dataset from Task 1D of the CLEF2022 CheckThat Lab Nakov et al. (2022). The SOTA number for this task was achieved by using Few-Shot GPT-4 (3-shot), as reported in Larabench Abdelali et al. (2024).
A.2 Check-worthiness
Check-worthiness helps streamline fact-checking by prioritizing claims most important for verification. This task operates as a binary classification, labeling tweets as either check-worthy or not check-worthy.
Dataset
For the Check-worthiness task, we utilized both the English and Arabic subsets of the dataset released for Task 1 of the CLEF2024 CheckThat Lab Barrón-Cedeño et al. (2024). The dataset includes tweets labeled with binary classifications: check-worthy or not check-worthy. The SOTA for this task was achieved using GPT-3.5 (fine-tuned) for Arabic and RoBERTa for English, as reported in Struß et al. (2024).
A.3 Claim
Claim identifies whether a piece of text contains a factual, verifiable statement. A factual claim is one that can be substantiated through reliable sources, such as statistics, reports, or witness accounts. This task is crucial for fact-checking systems, as it helps distinguish between statements that can be objectively verified and those that are subjective or opinion-based.
Dataset
For the Claim task, we used three binary-labeled datasets: two in Arabic and one in English. The English dataset, “claim-detection” is sourced from Nithiwat/claim-detection on Hugging Face Nithiwat (2022). One of the Arabic datasets, “CT22Claim” comes from the Arabic subset of Task 1B of the CLEF2022 CheckThat Lab Nakov et al. (2022). The second Arabic dataset, “ans-claim” consists of true and false claims generated using crowdsourcing, based on the Arabic News Texts (ANT) corpus Khouja (2020). The SOTA for the Arabic dataset “CT22Claim” was achieved using Zero-shot GPT-3.5 in Larabench Abdelali et al. (2024), while no SOTA was available for the English dataset “claim-detection.”
A.4 Cyberbullying
Cyberbullying identifies whether a piece of text contains abusive, harassing, or threatening behavior directed towards individuals online. This task plays a crucial role in moderating online spaces by flagging harmful content that can affect the well-being of users.
Dataset
For the Cyberbullying task, we used datasets in three languages: Arabic, English, and Hindi. The Arabic dataset, “ArCyc_CB” is sourced from Shannag (2023). The English dataset, “Cyberbullying” is developed by  Wang et al. (2020). The Hindi dataset, “MC-Hinglish1.0” is developed by  Laskar et al. (2024). The SOTA for ArCyc_CB was achieved using Support Vector Machine (SVM) with word embedding  Shannag et al. (2022). The English dataset “Cyberbullying” achieved its SOTA as reported inAlqahtani and Ilyas (2024), and for the Hindi dataset “MC-Hinglish1.0,” the SOTA was achieved using a voting classifier Laskar et al. (2024).
A.5 Emotion
Emotion focuses on determining whether a piece of text conveys an emotion and identifying which specific emotion is being expressed.
Dataset
For the Emotion task, we utilized two Arabic datasets and one English dataset. The Arabic datasets are “Emotional-Tone” Al-Khatib and El-Beltagy (2018) and “NewsHeadline” Galal Elsayed et al. (2022), while the English dataset is “emotion” Saravia et al. (2018). The SOTA for the Emotional-Tone dataset was achieved using the Naïve Bayes Algorithm with 10-fold cross-validation Al-Khatib and El-Beltagy (2018). The NewsHeadline dataset achieved 100% accuracy using Bag-of-Words (BOW) features Galal Elsayed et al. (2022). For the English dataset, the CARER model was used to achieve the SOTA Saravia et al. (2018).
A.6 Factuality
Factuality mainly focuses on assessing the truthfulness of a claim, determining whether the information presented is accurate or false.
Dataset
For the Factuality Detection task, we utilized five datasets: two in Arabic, two in English, and one in Hindi. The Arabic datasets are “Arafacts” Ali et al. (2021) and “COVID19Factuality” Alam et al. (2021b, a). The English datasets are “News_dataset” Dogo et al. (2020) which combines two datasets (true and fake news) into one. The second English Dataset is “Politifact” Misra (2022b). Finally, the Hindi dataset is “fake-news” Patil et al. (2024). The datasets focus on classifying claims as either true or false to assess their factuality. The SOTA for “COVID19Factuality” by Alam et al. (2021b, a). The SOTA for “News_dataset” was achieved using LSVM as a classifier Ahmed et al. (2017), and for “Politifact” by Rangapur et al. (2023). Moreover, Arafacts SOTA was obtained by  Othman et al. (2024). No SOTA was found for the fake-news in the Hindi dataset.
A.7 Harmful
Harmful Detection focuses on determining whether a piece of text contains harmful content, which may include rumors, offensive language, hate speech, cyberbullying, violence, as well as racist, misogynistic, or sexist remarks. This task is essential for curbing the spread of harmful information online. The approach is based on work proposed in Alam et al. (2021b); Nakov et al. (2022).
Dataset
For the Harmful Detection task, we utilized the Arabic dataset “CT22Harmful” Nakov et al. (2022), which is a binary dataset provided in Subtask 1C: Harmful tweet detection. The SOTA for this task was achieved byTaboubi et al. (2022).
A.8 Hate Speech
Hate Speech focuses on determining whether a piece of text contains hate speech. Hate speech refers to language that expresses hostility or animosity towards a specific group, or is intended to degrade, humiliate, or insult its members. This task is vital for monitoring online content and reducing the spread of harmful language.
Dataset
For the Hate Speech Detection task, we utilized four datasets: two in Arabic and two in Hindi. The Arabic datasets are “annotated-hatetweets-4-classes” Ahmad et al. (2024b), using the latest version 3, and “OSACT4SubtaskB” Mubarak et al. (2020a). The Hindi datasets are “hate-speech-detection” Das et al. (2022), renamed from its original name “HateCheckHIn”, and “Hindi-Hostility-Detection-CONSTRAINT-2021” Bhardwaj et al. (2020), which is a multi-labeled dataset. The SOTA for “annotated-hatetweets-4-classes” was achieved by Ahmad et al. (2024a), for “OSACT4SubtaskB” by Husain (2020), for “hate-speech-detection” by Das et al. (2022), and for “Hindi-Hostility-Detection-CONSTRAINT-2021” by Bhardwaj et al. (2020).
A.9 Natural Language Inference
Natural Language Inference focuses on identifying the relationship between two sentences. It involves determining whether the second sentence is logically supported by, contradicts, or remains neutral with respect to the first sentence.
Dataset
For the Natural Language Inference task, we used a binary-labeled Hindi dataset Gautam et al. (2021a), which combines two datasets released by Dhar et al. (2018) and Srivastava and Singh (2020). The SOTA for this task was achieved by Gautam et al. (2021b).
A.10 News Credibility
News Credibility involves determining whether a news article is reliable. This includes evaluating the article based on factors such as accuracy, fairness, objectivity, trustworthiness, completeness, and the presence or absence of biases.
Dataset
For the News Credibility task, we used one Arabic dataset: “NewsCredibilityDataset” Samdani et al. (2023).
A.11 News Genre Categorization
News Genre Categorization involves classifying news articles based on both their content and style. This task identifies the main topic or theme of an article, while also determining its genre, categorizing it as an opinion piece, objective news reporting, or satire.
Dataset
This task utilizes both Arabic and English datasets. The Arabic datasets include ASND” Chowdhury et al. (2020a), “SANADAkhbarona”, “SANADAlArabiya”, “SANADAlkhaleej” Einea et al. (2019), and “UltimateDataset” Al-Dulaimi (2022). The English datasets include “CNN_News_Articles_2011-2022” Hadasu (2022), “News_Category_Dataset” Misra (2022a); Misra and Grover (2021), with the latter containing 42 labels, and “SemEval23T3-subtask1” Piskorski et al. (2023). The SOTA for the SANAD datasets was achieved by Elnagar et al. (2020), for ASND by Chowdhury et al. (2020b), for UltimateDataset by Setu et al. (2024), for CNN_News_Articles_2011-2022 on Hugging Face by CHERGUELAINE Ayoub & BOUBEKRI Faycal, and for SemEval23T3-subtask1 by Piskorski et al. (2023).
A.12 Summarization
Summarization focuses on producing concise and coherent summaries of articles, capturing the key points in a clear and succinct manner.
Dataset
For the Summarization task, we used the "xlsum" multilingual dataset and the SOTA result  Hasan et al. (2021) across all three languages: English, Arabic, and Hindi.
A.13 Offensive Language
Offensive Language Detection identifies whether a piece of text contains offensive language. Offensive speech includes vulgar or targeted insults, explicit or implicit attacks against others, or the use of inappropriate language.
Dataset
Dataset for the Offensive Language Detection task, we used five datasets: two in Arabic, two in English, and one in Hindi. The Arabic datasets are “ArCyc _OFF” (Shannag, 2023) and “OSACT4SubtaskA” (Zampieri et al., 2020). The English datasets are “Offensive_Hateful_Dataset_New” (Christina, 2024) from Hugging Face, and “offensive_language_dataset” (Zampieri et al., 2019). The Hindi dataset is “Offensive Speech Detection” (Mathur et al., 2018). The labels for these datasets were extracted from Davidson et al. (2017). The SOTA for “ArCyc_OFF” was achieved by Shannag et al. (2022), for “Offensive Speech Detection” by Das et al. (2022), for “OSACT4SubtaskA” by Mubarak et al. (2020b), and for “offensive_language_dataset” by Dehghani (2024). The "hate-offensive-speech" dataset on Hugging Face achieved its SOTA by Purvesh Patel, while no SOTA was found for “Offensive_Hateful_Dataset_New”.
A.14 Propaganda
Propaganda detection focuses on identifying propaganda in a piece of text. Propaganda is a form of communication designed to influence people’s opinions or actions toward a specific goal, often using strategic rhetorical and psychological techniques.
Dataset
For the Propaganda task, we used two binary-labeled datasets: one in Arabic and one in English. The Arabic dataset and the SOTA which was achieved using AraBERT is “ArPro” Hasanain et al. (2024a), and the English dataset and SOTA is “QProp” Barrón-Cedeno et al. (2019).
A.15 Sarcasm
Sarcasm focuses on determining whether a piece of text conveys sarcasm or not.
Dataset
For the Sarcasm task, we used two binary-labeled datasets: one in Arabic and one in English. The Arabic dataset is “ArSarcasm-v2” Abu Farha et al. (2021), and the English dataset is “News-Headlines-Dataset-For-Sarcasm-Detection” Misra and Arora (2023); Misra and Grover (2021). The SOTA for both datasets is the same as the respective references where the datasets were found.
A.16 Sentiment
Sentiment classification involves in identifying and classifying sentiment expressed through text.
Dataset
For the Sentiment task, we used four datasets: two in Arabic, one in English, and one in Hindi. The Arabic dataset “ar_reviews_100k” Elnagar et al. (2018); Elnagar and Einea (2016) includes hotel and book reviews from the HARD and BRAD datasets, with additional airline reviews. The second Arabic dataset is “ArSAS” Elmadany et al. (2018). The English dataset is “NewsMTSC-dataset” Hamborg and Donnay (2021), and the Hindi dataset is “Sentiment Analysis” Joshi et al. (2016a). The SOTA for “NewsMTSC-dataset” was achieved by Hamborg and Donnay (2021), and for “Sentiment Analysis” by Joshi et al. (2016b). No SOTA was found for “ar_reviews_100k”.
A.17 Stance
Stance involves predicting the author’s position toward a particular subject based on a written text. The stance may be expressed explicitly or implied within the content.
Dataset
For the Stance task, we used two Arabic datasets: “Mawqif-Arabic-Stance-main” Alturayeif et al. (2022a) and “stance” Khouja (2020). The SOTA for “Mawqif-Arabic-Stance-main” was achieved by Alturayeif et al. (2022b), and for “stance”, the SOTA was achieved using pretraining (BERT), as reported in the same dataset reference Khouja (2020).
A.18 Subjectivity
Subjectivity involves determining whether a piece of text is subjective or objective. A sentence is considered subjective if it is influenced by personal feelings, tastes, or opinions; otherwise, it is classified as objective.
Dataset
For the Subjectivity task, we used two binary-labeled datasets: one in Arabic and one in English. The Arabic dataset is “ThatiAR” Suwaileh et al. (2024), and the English dataset is “CT24_T2” from the CLEF2024–CheckThat-Lab Task 2 Barrón-Cedeño et al. (2024). The SOTA for “ThatiAR” was achieved by Suwaileh et al. (2024), and the SOTA for “CT24_T2” was achieved by Gruman and Kosseim (2024).
Appendix B Dataset Sizes After Pre-processing
The Tables 3, 4 and 5 present the sizes of the datasets used in this study, after pre-processing. The datasets are categorized by language (Arabic, English, and Hindi) and we report the distribution of training, development test sets, along with the number of labels for each task.
Task Dataset # Labels # Train # Test # Dev
Attentionworthiness CT22Attentionworthy 9 2,470 1,186 1,071
Checkworthiness CT24_T1 2 22,403 1,031 318
Claim CT22Claim 2 3,513 1,248 339
Cyberbullying ArCyc_CB 2 3,145 900 451
Emotion Emotional-Tone 8 7,024 2,009 1,005
Emotion NewsHeadline 7 939 323 160
Factuality Arafacts 5 4,354 1,245 623
Factuality COVID19Factuality 2 3,513 988 339
Harmful CT22Harmful 2 2,484 1,201 1,076
Hate Speech annotated-hatetweets-4-classes 4 210,526 100,565 90,544
Hate Speech OSACT4SubtaskB 2 4,778 1,827 2,048
News Genre Categorisation ASND 10 74,496 21,942 11,136
News Genre Categorisation SANADAkhbarona 7 62,210 7,824 7,824
News Genre Categorisation SANADAlArabiya 6 56,967 7,123 7,120
News Genre Categorisation SANADAlkhaleej 7 36,391 4,550 4,550
News Genre Categorisation UltimateDataset 10 133,036 38,456 19,269
News Credibility NewsCredibilityDataset 2 8,671 2,730 1,426
Summarization xlsum 37,425 4,689 4,689
Offensive Language ArCyc_OFF 2 3,138 900 450
Offensive Language OSACT4SubtaskA 2 4,780 1,827 2,047
Propaganda ArPro 2 6,002 1,326 672
Sarcasm ArSarcasm-v2 2 8,749 2,996 3,761
Sentiment ar_reviews_100k 3 69,998 20,000 10,000
Sentiment ArSAS 4 13,883 3,976 1,987
Stance Mawqif-Arabic-Stance-main 2 3,162 560 950
Stance stance 3 2,652 379 755
Subjectivity ThatiAR 2 2,446 748 467
Table 3: Data distribution across Arabic datasets.
Task Dataset # Labels # Train # Test # Dev
Checkworthiness CT24_T1 2 22,403 1,031 318
Claim claim-detection 2 23,224 7,267 5,815
Cyberbullying Cyberbullying 6 32,551 9,473 4,751
Emotion emotion 6 280,551 82,454 41,429
Factuality News_dataset 2 28,147 8,616 4,376
Factuality Politifact 6 14,799 4,230 2,116
News Genre Categorisation CNN_News_Articles_2011-2022 6 32,193 5,682 9,663
News Genre Categorisation News_Category_Dataset 42 145,748 41,740 20,899
News Genre Categorization SemEval23T3-subtask1 3 302 83 130
Summarization xlsum 306,493 11,535 11,535
Offensive Language Offensive_Hateful_Dataset_New 2 42,000 5,252 5,254
Offensive Language offensive_language_dataset 2 29,216 3,653 3,653
Offensive/Hate-Speech hate-offensive-speech 3 48,944 2,799 2,802
Propaganda QProp 2 35,986 10,159 5,125
Sarcasm News-Headlines-Dataset-For-Sarcasm-Detection 2 19,965 5,719 2,858
Sentiment NewsMTSC-dataset 3 7,739 747 320
Subjectivity CT24_T2 2 825 484 219
Table 4: Data distribution across English datasets.
Task Dataset # Labels # Train # Test # Dev
Cyberbullying MC-Hinglish1.0 7 7,400 1,000 2,119
Factuality fake-news 2 8,393 2,743 1,417
Hate Speech hate-speech-detection 2 3,327 951 476
Hate Speech Hindi-Hostility-Detection-CONSTRAINT-2021 15 5,718 1,651 811
Natural Language Inference Natural Language Inference 2 1,251 447 537
Summarization xlsum 70,754 8,847 8,847
Offensive Speech Offensive Speech Detection 3 2,172 636 318
Sentiment Sentiment Analysis 3 10,039 1,259 1,258
Table 5: Data distribution across Hindi datasets.
Appendix C Instructions Generation
For each task, each dataset and each language, we use two effective closed models, GPT-4o and Claude-3.5 Sonnet, to generate instructions. These instructions will be used to create the final instruction dataset for LLM fine-tuning. We prompt the models to generate these instructions using the prompt in Table 6. For all generated instructions, we append the following suffix to further instruct the LLM to limit its responses to the labels/summary, to simplify post-processing at inference time: Return only the label without any explanation, justification or additional text. Table 7 shows examples of the generated instructions. Note that we only generate instructions for the user role, while we keep the system role fixed to that presented in Table 7.
Role Prompt
System You are an expert LLM developer with expertise in writing instructions to instruction-tune LLMs for users‘ tasks.
User We are creating an English instruction-following dataset for a/an [language] dataset called: [dataset name] covering the task of [task]. The user defined the task as follows: [task definition]. For that task, the labels include: [labels]. Write 10 very diverse and concise English instructions. Return the instructions as strings in a list format as follows [].
Table 6: Prompts used to generate instructions through LLMs.
Model Instruction System Role
GPT-4o Classify the given text as either ‘offensive’ or ‘not-offensive-hateful’. Return only the label without any explanation, justification or additional text. You are a social media expert providing accurate analysis and insights.
Claude-3.5 Evaluate whether the given text is ‘offensive’ or ‘not-offensive-hateful’, considering vulgar or targeted attacks. Return only the label without any explanation, justification or additional text. You are a social media expert providing accurate analysis and insights.
Table 7: Examples of instructions generated by two LLMs for the offensive language detection task on the English offensive_language_dataset, along with the pre-defined system role prompt.
Appendix D Data Release and License
The LlamaLens model and the instruction-following dataset will be publicly released under the Creative Commons Attribution Non Commercial Share Alike 4.0: https://creativecommons.org/licenses/by-nc-sa/4.0/.
Appendix E Use of AI assistant
We used AI assistants such as GPT-4o and Claude for generating the instructions dataset, as well as for spelling and grammar checking for the text of the paper.